# A module-theoretic interpretation of quantum expansion formula

### v1

@@ 4.2. Quantum Laurent expansions
Fix $s$. Assume that $a_{1_{s}}$$a_{4_{s}}$$\tau_{i_{s}}$ and $a_{2_{s}}$$a_{3_{s}}$$\tau_{i_{s}}$ are triangles in $\Gamma$ such that $a_{1_{s}}$$a_{3_{s}}$ are clockwise to $\tau_{i_{s}}$ and $a_{2_{s}}$$a_{4_{s}}$ are counterclockwise to $\tau_{i_{s}}$. Note that $a_{i_{s}}$, $i\in 1,2,3,4$ may be boundary arcs.
@@
Suppose that $P$ can twist on $G(p_{s})$. If the edges labeled $a_{2_{s}}$$a_{4_{s}}$ of $G(p_{s})$ are in $P$, define
@@
There is a bijection $\gamma\leftrightarrow X_{\gamma}$ between the internal arcs in $(S,M)$ and the cluster variable in $\mathcal{A}$. In particular, the arcs in $\Gamma$ are in one-to-one correspondence with the cluster variables in the initial seed. That is, $X_{\tau_{i}}=X^{e_{i}}$.
@@
The cluster monomial $x(P)$ associated to $P$ is
@@
[…]
@@ 4.3. Snake graphs and string modules
The theory of snake graphs is closely related to string modules(see [5],[22]). Here we recall some results in their paper. More specifically, there exists a bijection between string modules and snake graphs, up to a certain equivalence.
  • (From strings to snake graphs)Given a string $w$. If $w$ is a vertex then the corresponding snake graph is a single tile. Otherwise suppose $w=a_{1}\cdots a_{n}$ with $a_{i}\in\left\{\rightarrow,\leftarrow\right\}$.
@@
    Then for each vertex $x_{i}$ there is a tile $G(i)$ with diagonal $x_{i}$. If $a_{1}=\rightarrow$, then $G(2)$ is to the right of $G(1)$; otherwise, it is above $G(1)$. If $a_{i+1}$ and $a_{i}$ have the same direction, then $G(i),G(i+1),G(i+2)$ are zigzag. Otherwise, they are straight. Gluing all the $G(i)$ we get the snake graph $G(w)$.
  • (From snake graphs to strings)For any snake graph $G$ with tiles $G(1),\cdots,G(s)$, the string $w$ are concatenation of $a_{i}$ connecting the diagonals of $G(i)$ and $G(i+1)$. If $G(2)$ is to the right of $G(1)$, then $a_{1}=G(1)\rightarrow G(2)$. Otherwise, $a_{1}=G(1)\leftarrow G(2)$. If $G(i),G(i+1),G(i+2)$ are zigzag, then $a_{i}$ and $a_{i+1}$ have the same direction. Otherwise, they are inverse to each other.
In addition, if there exists an arc $\gamma$ in $(S,M)$ whose string is $w$, the snake graph $G(w)$ and $G_{\Gamma,\gamma}$ coincides. Moreover, this construction induces a bijection between the canonical submodules of $M(w)$ and perfect matchings of $G(w)$. More specifically, any perfect matching $P$ determines an index set in the following way: the symmetric difference $(P_{-}(G_{w})\cup P)\backslash(P_{-}(G_{w})\cap P)$ form a set of enclosed tiles of $G(w)$. The index set corresponding to these tiles will be the index set of the canonical submodule $M(P)$.
@@
where $e=(e_{1},e_{2},\dots,e_{n})$, and $e_{i}$ is the number of tiles in the symmetric difference whose diagonal is labeled $\tau_{i}$. Also $e$ is the dimension vector of $M(P)$. If $\gamma$ is the arc with string $w$, then the $g$-vector of $\gamma$ is given by
@@
Denote $x^{\Gamma}(P)=x^{ind_{\Gamma}(M(w))+B_{\Gamma}dimP}$ and $X^{\Gamma}(P)$ be the corresponding element in quantum cluster algebra, then
@@
and a quantum version $X(P)=X^{\Gamma}(P)$.
@@ 5. MODULE INTERPRETATION
In this section, we provide a module interpretation of the map $v$ in 4.16. The key observation is the following
@@
  • The number of edges label $\tau_{k}$ corresponds to the index of canonical submodules.
@@
Denote the vertices of $w$ by $x_{1},x_{2},\cdots,x_{s+1}$. Let $P$ be a perfect matching of $G(w)$. The minimal perfect matching of $G(w)$ is denoted by $P_{-}$. The symmetric difference provides a collection of enclosed tiles, which can give an index subset of $\left\{1,2\dots,s\right\}$. Under the equivalence above, we still denote the index set of the canonical submodule $M(P)$ by $P$. First we need to provide an algorithm of the number of edges.
Let $\tau_{k}$ be a vertex of $Q$. Then after choosing a suitable label the subquiver of neighborhood of $\tau_{k}$ and the triangle in surface are showed in Figure 1. Note that some vertices may be frozen.
@@
If there exists some $j$ with $2\leq j\leq s-1$ satisfying $x_{j}=\tau_{k}$ and $a_{j-1}$ and $a_{j}$ both are inverse, define two quantities to count the number of edges labeled $\tau_{k}$ in $E(G(j+1))\cap P$ and $E(G(j-1))\cap P$.
$n^{+}(\tau_{k},j,P)=\left\{\begin{aligned} 1&&&if\ j+1\in P\\ 0&&&otherwise.\end{aligned}\right$.
@@
The quantities defined here are use to count the number of edges in an equivalence class defined in [11]. There the edges incident to same diagonal who in addition all have the same label are equivalent. So the number in this equivalence class is $n^{+}+n^{-}$. The procedure requiring the count of rightward or leftward steps corresponds exactly to the definition of $G^{+}_{s}$ or $G^{-}_{s}$, respectively.
@@
If there exists some $j$ with $2\leq j\leq s-1$ satisfying $x_{j}=\tau_{k}$ and $a_{j-1}$ and $a_{j}$ both are direct, define two quantities to count the number of edges labeled $\tau_{k}$ in $E(G(j+1))\cap P$ and $E(G(j-1))\cap P$.
@@
Without loss of generality, assume $x_{j-1}=\tau_{k-1}$, $x_{j+1}=\tau_{k+1}$. The snake graph of G(j1),G(j),G(j+1)G(j-1),G(j),G(j+1) is showed in Figure 4. The blue edge always belongs to $P_{-}$ and $n^{+}$ just depends on whether the blue edge belong to $P$. According to the definition of symmetric difference, the blue edge is in $P$ if and only if $j+1\notin P$. And the description for $n^{-}$ is true for the same reason. ∎
@@
Without loss of generality, assume $x_{j-1}=\tau_{k-2}$, $x_{j+1}=\tau_{k+2}$. The snake graph of G(j1),G(j),G(j+1)G(j-1),G(j),G(j+1) is showed in Figure 5. The blue edges always belong to $P_{-}$ and the red edges don’t. $n^{+}$ just depends on whether the red edge belongs to $P$. According to the definition of symmetric difference, the red edge is in $P$ if and only if $j+1\in P$. And the description for $n^{-}$ is true for the same reason. ∎
@@
If there exists some $j$ satisfying $\left\{x_{j},x_{j+1}\right\}=\left\{\tau_{k-1},\tau_{k-2}\right\}$ or $\left\{\tau_{k+1},\tau_{k+2}\right\}$, let $n(\tau_{k},j,P)=1$ if exactly one of $j$ and $j+1$ belongs to $P$, otherwise $n(\tau_{k},j,P)=0$. Then $n(\tau_{k},j,P)$ count the number of edges in $P$ labeled $\tau_{k}$ in this two tiles.
@@
Without loss of generality, assume $x_{j}=\tau_{k-1}$, $x_{j+1}=\tau_{k-2}$. The snake graph of G(j1),G(j),G(j+1)G(j-1),G(j),G(j+1) is showed in Figure 6. The blue edges always don’t belong to $P_{-}$. $n(\tau_{k},j,P)$ just depends on whether the blue edge belongs to $P$. According to the definition of symmetric difference, the blue edge is in $P$ if and only exactly one of $j$ and $j+1$ belongs to $P$. ∎
@@
In the following, we need to consider the first or last tile and count the number of edges label $\tau_{k}$ in these tiles. The strategy is totally the same as above. We show some of example in Figure 7. For example, if the first tile has diagonal $\tau_{k-1}$ and it cannot be any case above, then the blue edge label $\tau_{k}$ belongs to $P_{-}$. So whether it belongs to $P$ is equivalent to whether it $1$ belongs to $P$. Similarly for the other case. We provide a summary of all results in this case here.
@@
If $x_{1}\in\left\{\tau_{k-1},\tau_{k+1}\right\}$ and $x_{2}\notin\left\{\tau_{k-2},\tau_{k-1},\tau_{k},\tau_{k+1},\tau_{k+2}\right\}$, then let $n(\tau_{k},1,P)=0$ if $1\in P$ and $n(\tau_{k},1,P)=1$ if $1\notin P$. If $x_{1}\in\left\{\tau_{k-2},\tau_{k+2}\right\}$, then $n(\tau_{k},1,P)=1$ if $1\in P$ and $n(\tau_{k},1,P)=0$ if $1\notin P$. In both case, $n(\tau_{k},1,P)$ counts the number of edges labeled $\tau_{k}$ in $P$ in the first tile.
And dually for the last tile
@@
If $x_{s}\in\left\{\tau_{k-1},\tau_{k+1}\right\}$ and $x_{s-1}\notin\left\{\tau_{k-2},\tau_{k-1},\tau_{k},\tau_{k+1},\tau_{k+2}\right\}$, then $n(\tau_{k},s,P)=1$ if $s\notin P$ and $n(\tau_{k},s,P)=0$ if $s\in P$. If $x_{s}\in\left\{\tau_{k-2},\tau_{k+2}\right\}$, then $n(\tau_{k},s,P)=0$ if $s\notin P$ and $n(\tau_{k},s,P)=1$ if $s\in P$. In both case, $n(\tau_{k},s,P)$ counts the number of edges labeled $\tau_{k}$ in $P$ in the last tile.
@@
In the last case, we need to consider the first or last two tile and count the number of edges label $\tau_{k}$ in these tiles. We show some of example in Figure 8. For example, if the first tile has diagonal $\tau_{k}$ and the second tile has diagonal $\tau_{k-2}$, then the blue edge belongs to $P_{-}$ and the red edge can’t belong to $P_{-}$. So whether the red edge belongs to $P$ is equivalent to whether it $2$ belongs to $P$. Similarly for the other case. We provide a summary of all results in this case here.
@@
If $x_{1}=\tau_{k}$ and $x_{2}\in\left\{\tau_{k-2},\tau_{k+2}\right\}$, then $n^{+}(\tau_{k},1,P)=1$ if $2\in P$ and $n^{+}(\tau_{k},1,P)=0$ if $2\notin P$. If $x_{2}\in\left\{\tau_{k-1},\tau_{k+1}\right\}$, then $n^{+}(\tau_{k},1,P)=1$ if $2\notin P$ and $n^{+}(\tau_{k},1,P)=0$ if $2\in P$. In both case, $n(\tau_{k},1,P)$ counts the number of edges labeled $\tau_{k}$ in $P$ in the first two tile.
And dually for the last two tile
@@
If $x_{s}=\tau_{k}$ and $x_{s-1}\in\left\{\tau_{k-2},\tau_{k+2}\right\}$, then $n^{-}(\tau_{k},s,P)=1$ if $s-1\in P$ and $n^{-}(\tau_{k},s,P)=0$ if $s-1\notin P$. If $x_{s-1}\in\left\{\tau_{k-1},\tau_{k+1}\right\}$, then $n^{-}(\tau_{k},s,P)=0$ if $s-1\in P$ and $n^{-}(\tau_{k},s,P)=1$ if $s-1\notin P$. In both case, $n(\tau_{k},s,P)$ counts the number of edges labeled $\tau_{k}$ in $P$ in the last two tile.
@@
    For the cases where $n(\tau_{k},j,P)$ does not present, let $n(\tau_{k},j,P)=n^{+}(\tau_{k},j,P)+n^{-}(\tau_{k},j,P)$.
@@
    In [11], the equivalence class of edges label $\tau_{k}$ must be one of the above case. So $n(\tau_{k},j,P)$ is the number of edges belonging to $P$ in the equivalence class near $x_{j}$.
@@
Here we define the number identical to Definition 4.9
@@
    If two submodules $N$ and $N^{\prime}$ with $N=N^{\prime}\cup{x_{j}}$ for an index $j$, then
@@
The quantum cluster variable asssociated with $M(w)$ can be expressed by
@@
The arcs in surface or their string modules are in bijection with the quantum cluster variables. Note that in Definition 4.10, the edges labeled $a_{2_{s}}$ and $a_{4_{s}}$ belonging to $P$ are equivalent to $p_{s}\in P$. So $\Omega^{\prime}(x_{j},P)=\Omega(j,P)$ and the map $v_{\gamma}$ has the same initial condition and recursion formula. Therefore, the map $v_{\gamma}$ is equal to the map $v$ in Theorem 4.16. For any $P$, we also have $X(P)=X^{\Gamma}(P)$. Then the formula is true by the bijection for perfect matchings and canonical submodules. ∎

### v2

@@ 4.2. Quantum Laurent expansions
Fix $s$. Assume that $a_{1_{s}}$, $a_{4_{s}}$, $\tau_{i_{s}}$ and $a_{2_{s}}$, $a_{3_{s}}$, $\tau_{i_{s}}$ are triangles in $\Gamma$ such that $a_{1_{s}}$, $a_{3_{s}}$ are clockwise to $\tau_{i_{s}}$ and $a_{2_{s}}$, $a_{4_{s}}$ are counterclockwise to $\tau_{i_{s}}$. Note that $a_{i_{s}}$, for $i\in 1,2,3,4$ may be boundary arcs.
@@
Suppose that $P$ can twist on $G(p_{s})$. If the edges labeled $a_{2_{s}}$, $a_{4_{s}}$ of $G(p_{s})$ are in $P$, define
@@
There is a bijection $\gamma\leftrightarrow X_{\gamma}$ between the internal arcs in $(S,M)$ and the cluster variables in $\mathcal{A}$. In particular, the arcs in $\Gamma$ are in one-to-one correspondence with the cluster variables in the initial seed. That is, $X_{\tau_{i}}=X^{e_{i}}$.
@@
The cluster monomial $x(P)$ associated with $P$ is
@@
    (initial conditions) $v(P_{-}(G_{\Gamma,\gamma}))=v(P_{+}(G_{\Gamma,\gamma}))=0$
@@ 4.3. Snake graphs and string modules
The theory of snake graphs is closely related to string modules (see [6], [11]). Here we recall some results in their papers. More specifically, there exists a bijection between string modules and snake graphs, up to a certain equivalence.
  • (From strings to snake graphs) Given a string $w$. If $w$ is a vertex, then the corresponding snake graph is a single tile. Otherwise, suppose $w=a_{1}\cdots a_{n}$ with $a_{i}\in\left\{\rightarrow,\leftarrow\right\}$.
@@
    Then, for each vertex $x_{i}$, there is a tile $G(i)$ with diagonal $x_{i}$. If $a_{1}=\rightarrow$, then $G(2)$ is to the right of $G(1)$; otherwise, it is above $G(1)$. If $a_{i+1}$ and $a_{i}$ have the same direction, then $G(i),G(i+1),G(i+2)$ are zigzag. Otherwise, they are straight. Gluing all the $G(i)$, we get the snake graph $G(w)$.
  • (From snake graphs to strings) For any snake graph $G$ with tiles $G(1),\cdots,G(s)$, the string $w$ is the concatenation of the $a_{i}$ connecting the diagonals of $G(i)$ and $G(i+1)$. If $G(2)$ is to the right of $G(1)$, then $a_{1}=G(1)\rightarrow G(2)$. Otherwise, $a_{1}=G(1)\leftarrow G(2)$. If $G(i),G(i+1),G(i+2)$ are zigzag, then $a_{i}$ and $a_{i+1}$ have the same direction. Otherwise, they are inverses to each other.
In addition, if there exists an arc $\gamma$ in $(S,M)$ whose string is $w$, the snake graph $G(w)$ and $G_{\Gamma,\gamma}$ coincide. Moreover, this construction induces a bijection between the canonical submodules of $M(w)$ and perfect matchings of $G(w)$. More specifically, any perfect matching $P$ determines an index set in the following way: the symmetric difference $(P_{-}(G_{w})\cup P)\backslash(P_{-}(G_{w})\cap P)$ forms a set of enclosed tiles of $G(w)$. The index set corresponding to these tiles is the index set of the canonical submodule $M(P)$.
@@
where $e=(e_{1},e_{2},\dots,e_{n})$, and $e_{i}$ is the number of tiles in the symmetric difference whose diagonal is labeled $\tau_{i}$. Also, $e$ is the dimension vector of $M(P)$. If $\gamma$ is the arc with string $w$, then the $g$-vector of $\gamma$ is given by
@@
Denote $x^{\Gamma}(P)=x^{ind_{\Gamma}(M(w))+B_{\Gamma}dimP}$, and let $X^{\Gamma}(P)$ be the corresponding element in the quantum cluster algebra, then
@@
and the quantum version $X(P)=X^{\Gamma}(P)$.
@@ 5. MODULE INTERPRETATION
In this section, we provide a module interpretation of the map $v$ in 4.16. The key observation is the following:
@@
  • The number of edges labeled $\tau_{k}$ corresponds to the index of canonical submodules.
@@
Denote the vertices of $w$ by $x_{1},x_{2},\cdots,x_{s+1}$. Let $P$ be a perfect matching of $G(w)$. The minimal perfect matching of $G(w)$ is denoted by $P_{-}$. The symmetric difference provides a collection of enclosed tiles, which can give an index subset of $\left\{1,2,\dots,s\right\}$. Under the equivalence above, we still denote the index set of the canonical submodule $M(P)$ by $P$. First we need to provide an algorithm for the number of edges.
Let $\tau_{k}$ be a vertex of $Q$. Then after choosing a suitable label the subquiver of the neighborhood of $\tau_{k}$ and the triangle in the surface are shown in Figure 1. Note that some vertices may be frozen.
@@
If there exists some $j$ with $2\leq j\leq s-1$ satisfying $x_{j}=\tau_{k}$ and $a_{j-1}$ and $a_{j}$ are both inverse, define two quantities to count the number of edges labeled $\tau_{k}$ in $E(G(j+1))\cap P$ and $E(G(j-1))\cap P$.
$n^{+}(\tau_{k},j,P)=\left\{\begin{aligned} 1&&&if\ j+1\in P\\ 0&&&otherwise\end{aligned}\right$.
@@
The quantities defined here are used to count the number of edges in an equivalence class defined in [18]. There the edges incident to the same diagonal which in addition all have the same label are equivalent. So the number in this equivalence class is $n^{+}+n^{-}$. The procedure requiring the count of rightward or leftward steps corresponds exactly to the definition of $G^{+}_{s}$ or $G^{-}_{s}$, respectively.
@@
If there exists some $j$ with $2\leq j\leq s-1$ satisfying $x_{j}=\tau_{k}$ and $a_{j-1}$ and $a_{j}$ are both direct, define two quantities to count the number of edges labeled $\tau_{k}$ in $E(G(j+1))\cap P$ and $E(G(j-1))\cap P$.
@@
Without loss of generality, assume $x_{j-1}=\tau_{k-1}$, $x_{j+1}=\tau_{k+1}$. The snake graph of G(j1),G(j),G(j+1)G(j-1),G(j),G(j+1) is shown in Figure 4. The red edge always belongs to $P_{-}$ and $n^{+}$ just depends on whether the blue edge belongs to $P$. According to the definition of symmetric difference, the blue edge is in $P$ if and only if $j+1\notin P$. The description for $n^{-}$ is true for the same reason. ∎
@@
Without loss of generality, assume $x_{j-1}=\tau_{k-2}$, $x_{j+1}=\tau_{k+2}$. The snake graph of G(j1),G(j),G(j+1)G(j-1),G(j),G(j+1) is shown in Figure 5. The blue edges always belong to $P_{-}$ and the red edges do not. $n^{+}$ just depends on whether the red edge belongs to $P$. According to the definition of symmetric difference, the red edge is in $P$ if and only if $j+1\in P$. The description for $n^{-}$ is true for the same reason. ∎
@@
If there exists some $j$ satisfying $\left\{x_{j},x_{j+1}\right\}=\left\{\tau_{k-1},\tau_{k-2}\right\}$ or $\left\{\tau_{k+1},\tau_{k+2}\right\}$, let $n(\tau_{k},j,P)=1$ if exactly one of $j$ and $j+1$ belongs to $P$, otherwise $n(\tau_{k},j,P)=0$. Then $n(\tau_{k},j,P)$ counts the number of edges in $P$ labeled $\tau_{k}$ in these two tiles.
@@
Without loss of generality, assume $x_{j}=\tau_{k-1}$, $x_{j+1}=\tau_{k-2}$. The snake graph of G(j1),G(j),G(j+1)G(j-1),G(j),G(j+1) is shown in Figure 6. The blue edges do not belong to $P_{-}$. $n(\tau_{k},j,P)$ just depends on whether the blue edge belongs to $P$. According to the definition of symmetric difference, the blue edge is in $P$ if and only if exactly one of $j$ and $j+1$ belongs to $P$. ∎
@@
In the following, we need to consider the first or last tile and count the number of edges labeled $\tau_{k}$ in these tiles. The strategy is totally the same as above. We show some examples in Figure 7. For example, if the first tile has diagonal $\tau_{k-1}$ and it cannot be any case above, then the blue edge labeled $\tau_{k}$ belongs to $P_{-}$. So whether it belongs to $P$ is equivalent to whether $1$ belongs to $P$. Similarly for the other case. We provide a summary of all results in this case here.
@@
If $x_{1}\in\left\{\tau_{k-1},\tau_{k+1}\right\}$ and $x_{2}\notin\left\{\tau_{k-2},\tau_{k-1},\tau_{k},\tau_{k+1},\tau_{k+2}\right\}$, then let $n(\tau_{k},1,P)=0$ if $1\in P$ and $n(\tau_{k},1,P)=1$ if $1\notin P$. If $x_{1}\in\left\{\tau_{k-2},\tau_{k+2}\right\}$, then $n(\tau_{k},1,P)=1$ if $1\in P$ and $n(\tau_{k},1,P)=0$ if $1\notin P$. In both cases, $n(\tau_{k},1,P)$ counts the number of edges labeled $\tau_{k}$ in $P$ in the first tile.
Dually, for the last tile
@@
If $x_{s}\in\left\{\tau_{k-1},\tau_{k+1}\right\}$ and $x_{s-1}\notin\left\{\tau_{k-2},\tau_{k-1},\tau_{k},\tau_{k+1},\tau_{k+2}\right\}$, then $n(\tau_{k},s,P)=1$ if $s\notin P$ and $n(\tau_{k},s,P)=0$ if $s\in P$. If $x_{s}\in\left\{\tau_{k-2},\tau_{k+2}\right\}$, then $n(\tau_{k},s,P)=0$ if $s\notin P$ and $n(\tau_{k},s,P)=1$ if $s\in P$. In both cases, $n(\tau_{k},s,P)$ counts the number of edges labeled $\tau_{k}$ in $P$ in the last tile.
@@
In the last case, we need to consider the first or last two tiles and count the number of edges labeled $\tau_{k}$ in these tiles. We show some examples in Figure 8. For example, if the first tile has diagonal $\tau_{k}$ and the second tile has diagonal $\tau_{k-2}$, then the blue edge belongs to $P_{-}$ and the red edge cannot belong to $P_{-}$. So whether the red edge belongs to $P$ is equivalent to whether $2$ belongs to $P$. Similarly for the other case. We provide a summary of all results in this case here.
@@
If $x_{1}=\tau_{k}$ and $x_{2}\in\left\{\tau_{k-2},\tau_{k+2}\right\}$, then $n^{+}(\tau_{k},1,P)=1$ if $2\in P$ and $n^{+}(\tau_{k},1,P)=0$ if $2\notin P$. If $x_{2}\in\left\{\tau_{k-1},\tau_{k+1}\right\}$, then $n^{+}(\tau_{k},1,P)=1$ if $2\notin P$ and $n^{+}(\tau_{k},1,P)=0$ if $2\in P$. In both cases, $n(\tau_{k},1,P)$ counts the number of edges labeled $\tau_{k}$ in $P$ in the first two tiles.
Dually, for the last two tile
@@
If $x_{s}=\tau_{k}$ and $x_{s-1}\in\left\{\tau_{k-2},\tau_{k+2}\right\}$, then $n^{-}(\tau_{k},s,P)=1$ if $s-1\in P$ and $n^{-}(\tau_{k},s,P)=0$ if $s-1\notin P$. If $x_{s-1}\in\left\{\tau_{k-1},\tau_{k+1}\right\}$, then $n^{-}(\tau_{k},s,P)=0$ if $s-1\in P$ and $n^{-}(\tau_{k},s,P)=1$ if $s-1\notin P$. In both cases, $n(\tau_{k},s,P)$ counts the number of edges labeled $\tau_{k}$ in $P$ in the last two tiles.
@@
    For the cases where $n(\tau_{k},j,P)$ is not present, let $n(\tau_{k},j,P)=n^{+}(\tau_{k},j,P)+n^{-}(\tau_{k},j,P)$.
@@
    In [18], the equivalence class of edges labeled $\tau_{k}$ must be one of the above cases. So $n(\tau_{k},j,P)$ is the number of edges belonging to $P$ in the equivalence class near $x_{j}$.
@@
Here we define numbers identical to those in Definition 4.9.
@@
    If two submodules $N$ and $N^{\prime}$ satisfy $N=N^{\prime}\cup{x_{j}}$ for an index $j$, then
@@
The quantum cluster variable associated with $M(w)$ can be expressed by
@@
The arcs in the surface or their string modules are in bijection with the quantum cluster variables. Note that in Definition 4.10, the edges labeled $a_{2_{s}}$ and $a_{4_{s}}$ belonging to $P$ are equivalent to $p_{s}\in P$. So $\Omega^{\prime}(x_{j},P)=\Omega(j,P)$ and the map $v_{\gamma}$ has the same initial condition and recursion formula. Therefore, the map $v_{\gamma}$ is equal to the map $v$ in Theorem 4.16. For any $P$, we also have $X(P)=X^{\Gamma}(P)$. Then the formula is true by the bijection between perfect matchings and canonical submodules. ∎